OpenStudy (anonymous):

[10.02] Solve the proportion 12 over 30 equals x over 5.. What is the value of x? 2 3 4 12

OpenStudy (goldphenoix):

\[\large \large \frac{ 12 }{ 30 } = \frac{ x }{ 5 }\] For any proportion, you would want to do cross multiply. So cross multiply: 30 * x = ? 12 * 5 = ?

OpenStudy (goldphenoix):

Or in this case you can do, 30 divide by what number would give us 5?

OpenStudy (anonymous):

6

OpenStudy (goldphenoix):

Okay. So 30 / 6 = 5. If you divide the denominator by 6, then you must divide the numerator by 6 as well; similar to an equation. So, 12 / 6 = ?

OpenStudy (anonymous):

2

OpenStudy (goldphenoix):

Yes. So x = 2. I hope this help! :)

OpenStudy (anonymous):

thank you